# On the Verification of Control Flow Attestation Evidence

### v2

@@ 7. Conclusion
We study if and how run-time evidence of CFA architectures can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify binary patches. We analyze and classify types of run-time evidence generated by current CFA architectures and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We introduce SABRE to verify path validity and to enable $\sf{\mathcal{V}rf}$ to locate/patch vulnerabilities using only the application binary and run-time evidence. We implement and evaluate the public prototype of SABRE (Caulfield et al., 2024b) to demonstrate its effectiveness and compatibility with hardware-based and TEE-based CFA architectures.

### v3

@@ 7. Conclusion
We study how CFA evidence can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify binary patches. We analyze and classify types of CFA evidence and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We introduce SABRE to verify path validity and to locate/patch vulnerabilities using only the application binary and CFA evidence. We implement and evaluate SABRE’s public prototype (Caulfield et al., 2025), which demonstrates its effectiveness atop hardware-based and TEE-based CFA architectures.
Acknowledgments. We thank the anonymous reviewers and shepherd for their guidance.
APPENDIX
Appendix A Alternative Prover Configurations
Alternative mechanisms such as Address Space Layout Randomization (ASLR) (Shacham et al., 2004) or CFA optimizations (e.g., SpecCFA (Caulfield et al., 2024a)) may be concurrently employed on $\sf{\mathcal{P}rv}$ for increased security/performance. In both cases, SABRE remains effective as long as $\sf{\mathcal{V}rf}$ knows the configuration of the respective mechanism. For example, for CFA in MCUs (with no MMU), ASLR would be performed before $\sf{\mathcal{P}rv}$ deployment. Therefore, $\sf{\mathcal{V}rf}$ would have access to the randomized binary. Since an ASLR bypass would corrupt a pointer to a memory address post-randomization, SABRE would not need to perform any additional handling and can pass the randomized binary with the corresponding E2 evidence to SABRE. CFA optimizations on E2 evidence, including speculation-based ones such as SpecCFA (Caulfield et al., 2024a), are lossless by design. In it, $\sf{\mathcal{V}rf}$ replaces expected subpaths with reduced-sized reserved symbols. Since the subpath-to-symbol correspondence is known by $\sf{\mathcal{V}rf}$, its ability to analyze received evidence is not lost, regardless of this optimization.
Appendix B Limitations and Future Work
Scaling to Complex Systems. Most current CFA and run-time auditing targets MCUs (Ammar et al., 2024). Conversely, SABRE considers MCU binaries (i.e., MSP430 and ARM Cortex-M33). An open problem in this space is scaling $\sf{\mathcal{P}rv}$ RoTs from low-end MCUs to complex systems. Assuming CFA and generation of E2 in complex systems is possible, a scalability challenge would come from performing SABRE’s symbolic data-flow analysis. Although SABRE does not face typical path explosion risk in symbolic execution due to its use of concrete run-time traces to guide the execution, a unique concern would arise when emulating the active state (memory and registers) during symbolic data flow analysis, which for MCU binaries is already the bottleneck in SABRE’s Exploit Locator (recall Sec. 5.1). Optimizing SABRE’s state modeling during symbolic data-flow analysis is an interesting avenue for future work.
Adding Vulnerabilities and Attacks. This work focuses on control flow hijacking and code-reuse attacks. For this reason, we tailored our framework towards the most prevalent root-cause memory vulnerabilities for these attacks: use-after-free vulnerabilities and buffer overflow-induced out-of-bounds writes ((2024), MITRE). Additionally, these memory vulnerabilities are prevalent for launching arbitrary code execution in real-world MCU software (as discussed in Sec. 2.2). However, other memory vulnerabilities (e.g., signedness errors, integer overflows, and double frees) could lead to arbitrary code execution. In these cases, SABRE would determine that a generated patch was ineffective and generate a report to provide a starting point for manual inspection (as discussed in Sec. 4.4). Similarly, SABRE does not support data-oriented attacks (e.g., direct-data-manipulation or data-oriented-programming (Cheng et al., 2019)) and would fail in the same manner in its attempts to generate a patch or pass E2 verification if the attack did not corrupt any control-data. These attacks could be detected through DFA run-time evidence that includes data inputs (recall Sec. 2.3). Future work could extend SABRE to detect/patch other memory vulnerabilities and to use data inputs from DFA to address data-oriented attacks.